OpenStudy (anonymous):

(y+6/y+3)=(y-2/y-1)+1

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

I would suggest multiplying both sides of the equation by the common denominator (y+3)(y-1) and see what you have after that.

myininaya (myininaya):

\[\frac{y+6}{y+3}(y-1)(y+3)=\frac{y-2}{y-1}(y-1)(y+3)+1(y-1)(y+3)\]

myininaya (myininaya):

(y+6)(y-1)=(y-2)(y+3)+(y-1)(y+3)

OpenStudy (anonymous):

I'd say they can take a stab at it from there myin.

myininaya (myininaya):

k polpak hi by the way

OpenStudy (anonymous):

=)

OpenStudy (anonymous):

Thank you to both of you! :-)

OpenStudy (anonymous):

I apologize but I am not sure where to go from the last post that myin put. Do I combine the (y+3)'s and make it 2(y+3)?

myininaya (myininaya):

\[y^2+5y-6=(y+3)[(y-2)+(y-1)]\] \[y^2+5y-6=(y+3)(2y-3)\]

OpenStudy (anonymous):

And this is for solving for y?

myininaya (myininaya):

isn't that what you wanted to do here? i assumed u wanted to solve for y

OpenStudy (anonymous):

Yes that is what I need to do for the problem but I am getting confused. I apologize.

OpenStudy (anonymous):

You foil out those products, combine like terms, move everything to one side (equals 0) and solve the quadratic.

myininaya (myininaya):

where?

OpenStudy (anonymous):

Ok. Let me see if I get the right answer. Thank you.

OpenStudy (anonymous):

If you have questions about a step feel free to ask.

OpenStudy (anonymous):

But try stuff before you ask and explain why you're stuck.

OpenStudy (anonymous):

should I get y=-1,3?